# CFTs with Large Gap from Barnes-Wall Lattice Orbifolds

### v1

@@ Acknowledgements
We thank Klaus Lux and Sven Möller for helpful discussions and comments on the draft. We thank Thomas Gemünden and Kristine Van for initial collaboration on this project. The work of CAK is supported in part by the Simons Foundation Grant No. 629215. The work of CAK and JR is supported by NSF Grant 2111748.
@@ A.2. $E(m)$ and its lifts
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\qquad\sigma_{2}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\ .$ (A.2.1)
@@
$[z,{\Sigma_{ar}}]=1\qquad z^{2}={\Sigma_{ar}}^{2}=1\qquad[{\Sigma_{ar}},{\Sigma_{bs}}]=z^{\delta_{a,b}\delta_{r,s+1\mod 2}}$ (A.2.3)
@@
${\Sigma_{11}}=\begin{pmatrix}-I&0&2I&0\\ 0&-I&0&I\\ 0&0&I&0\\ 0&0&0&I\end{pmatrix}\ ,\qquad{\Sigma_{12}}=\begin{pmatrix}I&0&0&0\\ 0&I&0&0\\ I&0&-I&0\\ 0&2I&0&-I\end{pmatrix}\ ,$ (A.2.7)
where $I$ and 00 are block matrices of size $2^{m-2}$. Similar expressions hold for all other ${\Sigma_{ar}}$ after different permutations of the basis vectors. For our purposes however we of course need explicit expressions for the same choice of basis for all generators.
@@
    Consider $w_{i}A_{a2}H_{1}$ and $w_{i}H_{1}A_{a2}^{T}$. Assume $\mathbb{p}[i]=0$ and $i[m-a]=1$, since otherwise both terms vanish individually. (Note that because $m$ is odd, $\neg$ flips parity.) We have $w_{i}A_{a2}H_{1}=w_{\neg(i-e_{m-a})}$ if $i[m-1]=1$ and 0 otherwise, since the digit $j[m-1]$ determines whether $j>\neg j$. Similarly we have $w_{i}H_{1}A_{a2}^{T}=w_{\neg i+e_{m-a}}=w_{\neg(i-e_{m-a}}$ again if $i[m-1]=1$ and 0 otherwise. It follows that $w_{i}A_{a2}H_{1}=w_{i}H_{1}A_{a2}^{T}$. A similar argument works for $A_{a1}$.
Note that this argument only works because $A_{ai}$ leaves $j[m-1]$ invariant for $a>1$. For $a=1$ we do indeed find a different outcome:
@@
    Consider $w_{i}A_{12}H_{1}$. Assume $i[m-1]=1$, since otherwise the identity is automatically satisfied. But then $w_{i}A_{12}=w_{j}$ where $j[m-1]=0$, so that $\neg j>j$ and therefore $w_{j}H_{1}=0$.
@@
    Consider $w_{i}H_{1}A_{11}^{T}$. Assume $i[m-1]=1$, since otherwise $\neg i>i$ and therefore $w_{i}H_{1}=0$. But then $w_{i}H_{1}=j$ where $j[m-1]=0$, so that $w_{j}A_{11}^{T}=0$.
@@ Appendix B Conjugacy classes of $E(m)$ in $BRW(m)$
Start with an element $g$ of the form above. If all the factors are $\mathbb{1}_{2}$, then we are immediately in the case ±$\pm\mathbb{1}$. Assume therefore that there is at least one non-trivial factor. Then use (1) until there is only one or no $\sigma_{3}$ left.
@@ C.1. Schur averaging
The remaining coordinates then factor into sums. Because there is only a linear term (since $y_{i}^{2}=y_{i}$), each sum is either 0 or 2, depending if $\tilde{\beta}^{h}_{i}$ is 0 or 1. More precisely, if we define $\pi_{d}^{h}$ to be the projector onto the last last $\dim K-2r$ entries and $\tilde{\beta}^{h}_{d}=\tilde{\beta}^{h}\pi^{h}_{d}$, then the sum vanishes unless
@@ D.1. Intertwiner representation for $m=5$
$T_{12}$ ::: ${\underline{\check{e}}}=\{3,4\}\ ,{\underline{\check{f}}}=0$ (D.1.4)
$T_{22}$ ::: ${\underline{\check{e}}}=\{10\}\ ,{\underline{\check{f}}}=0$ (D.1.5)
$T_{51}$ ::: ${\underline{\check{e}}}=\{3,4\}\ ,{\underline{\check{f}}}=0$ (D.1.6)
else ::: ${\underline{\check{e}}}=0\ ,{\underline{\check{f}}}=0$ (D.1.7)
@@
({}$\left(\begin{array}[]{cccccccccc}\{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{10\}&\{\}&\{10\}&\{\}&\{10\}&\{10\}&\{10\}&\{10\}&\{\}&\{10\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{10\}&\{\}&\{10\}&\{\}&\{3,4\}&\{10\}&\{10\}&\{10\}&\{\}&\{10\}\\ \{\}&\{10\}&\{\}&\{3,4\}&\{\}&\{\}&\{\}&\{\}&\{3,4,13,14\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{3,4\}&\{\}&\{3,4\}&\{\}&\{3,4,13,14\}&\{3,4\}&\{3,4\}&\{3,4\}&\{\}&\{3,4\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \end{array}\right)$ (D.1.9)

### v2

@@ Acknowledgements
We thank Klaus Lux and Sven Möller for helpful discussions and comments on the draft. We thank Thomas Gemünden and Kristine Van for initial collaboration on this project. We thank an anonymous referee for comments on the draft. The work of CAK is supported in part by the Simons Foundation Grant No. 629215. The work of CAK and JR is supported by NSF Grant 2111748.
@@ A.2. $E(m)$ and its lifts
$\sigma_{1}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\ ,\qquad\sigma_{2}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\$. (A.2.1)
@@
$z,{\Sigma_{ar}}]=1\qquad z^{2}={\Sigma_{ar}}^{2}=1\qquad[{\Sigma_{ar}},{\Sigma_{bs}}]=z^{\delta_{a,b}\delta_{r,s+1\mod 2}}$ (A.2.3)
@@
${\Sigma_{11}}=\begin{pmatrix}-I&0&2I&0\\ 0&-I&0&I\\ 0&0&I&0\\ 0&0&0&I\end{pmatrix}\ ,\qquad{\Sigma_{12}}=\begin{pmatrix}I&0&0&0\\ 0&I&0&0\\ I&0&-I&0\\ 0&2I&0&-I\end{pmatrix}\$, (A.2.7)
where $I$ and 0 are block matrices of size $2^{m-2}$. Similar expressions hold for all other $\Sigma_{ar}$} after different permutations of the basis vectors. For our purposes however we of course need explicit expressions for the same choice of basis for all generators.
@@
    Consider $w_{i}A_{a2}H_{1}$ and $w_{i}H_{1}A_{a2}^{T}$. Assume $\mathbb{p}[i]=0$ and i[ma]=1i[m-a]=1, since otherwise both terms vanish individually. (Note that because $m$ is odd, $\neg$ flips parity.) We have $w_{i}A_{a2}H_{1}=w_{\neg(i-e_{m-a})}$ if i[m1]=1i[m-1]=1 and 0 otherwise, since the digit j[m1]j[m-1] determines whether $j>\neg j$. Similarly we have $w_{i}H_{1}A_{a2}^{T}=w_{\neg i+e_{m-a}}=w_{\neg(i-e_{m-a}}$ again if i[m1]=1i[m-1]=1 and 0 otherwise. It follows that $w_{i}A_{a2}H_{1}=w_{i}H_{1}A_{a2}^{T}$. A similar argument works for $A_{a1}$.
Note that this argument only works because $A_{ai}$ leaves j[m1]j[m-1] invariant for $a>1$. For $a=1$ we do indeed find a different outcome:
@@
    Consider $w_{i}A_{12}H_{1}$. Assume i[m1]=1i[m-1]=1, since otherwise the identity is automatically satisfied. But then $w_{i}A_{12}=w_{j}$ where j[m1]=0j[m-1]=0, so that $\neg j>j$ and therefore $w_{j}H_{1}=0$.
@@
    Consider $w_{i}H_{1}A_{11}^{T}$. Assume i[m1]=1i[m-1]=1, since otherwise $\neg i>i$ and therefore $w_{i}H_{1}=0$. But then $w_{i}H_{1}=j$ where j[m1]=0j[m-1]=0, so that $w_{j}A_{11}^{T}=0$.
@@ Appendix B Conjugacy classes of $E(m)$ in $BRW(m)$
Start with an element $g$ of the form above. If all the factors are $\mathbb{1}_{2}$, then we are immediately in the case $\pm\mathbb{1}$. Assume therefore that there is at least one non-trivial factor. Then use (1) until there is only one or no $\sigma_{3}$ left.
@@ C.1. Schur averaging
The remaining coordinates then factor into sums. Because there is only a linear term (since $y_{i}^{2}=y_{i}$), each sum is either 0 or 2, depending if $\tilde{\beta}^{h}_{i}$ is 0 or 1. More precisely, if we define $\pi_{d}^{h}$ to be the projector onto the last $\dim K-2r$ entries and $\tilde{\beta}^{h}_{d}=\tilde{\beta}^{h}\pi^{h}_{d}$, then the sum vanishes unless
@@ D.1. Intertwiner representation for $m=5$
$T_{12}$ $:$ ${\underline{\check{e}}}=\{3,4\}\ ,{\underline{\check{f}}}=0$ (D.1.4)
$T_{22}$ $:$ ${\underline{\check{e}}}=\{10\}\ ,{\underline{\check{f}}}=0$ (D.1.5)
$T_{51}$ $:$ ${\underline{\check{e}}}=\{3,4\}\ ,{\underline{\check{f}}}=0$ (D.1.6)
else $:$ ${\underline{\check{e}}}=0\ ,{\underline{\check{f}}}=0$ (D.1.7)
@@
$\left(\begin{array}[]{cccccccccc}\{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{10\}&\{\}&\{10\}&\{\}&\{10\}&\{10\}&\{10\}&\{10\}&\{\}&\{10\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{10\}&\{\}&\{10\}&\{\}&\{3,4\}&\{10\}&\{10\}&\{10\}&\{\}&\{10\}\\ \{\}&\{10\}&\{\}&\{3,4\}&\{\}&\{\}&\{\}&\{\}&\{3,4,13,14\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \{3,4\}&\{\}&\{3,4\}&\{\}&\{3,4,13,14\}&\{3,4\}&\{3,4\}&\{3,4\}&\{\}&\{3,4\}\\ \{\}&\{10\}&\{\}&\{10\}&\{\}&\{\}&\{\}&\{\}&\{3,4\}&\{\}\\ \end{array}\right)$ (D.1.9)